OpenStudy (anonymous):

The circular bull’s-eye of an archery target has a diameter of 24 cm, which is surrounded by 4 concentric rings with a width of 12 cm each. Draw the target in a coordinate plane with center at the origin. Write the equations of the circles that form the boundaries of the different regions of the target.

OpenStudy (imstuck):

It appears that the bull's eye is the center of the target and it has a diameter of 24 cm and a center at (0,0). Do you know the standard form for the equation of a circle?

OpenStudy (anonymous):

\[(x-h)^{2}+(y-k)^{2}=r ^{2}\]

OpenStudy (imstuck):

That's right. In your case, what is the values for the h and the k?

OpenStudy (anonymous):

I'm not sure?

OpenStudy (imstuck):

Your problem tells you. The bull's eye is centered at the origin. The origin has coordinates of (0,0). So your h is a 0, and your k is a 0. That means that the center of the circle is right at the place where the x and y axes cross. So far your equation is this: x^2 + y^2 =

OpenStudy (imstuck):

What is the radius of the bull's eye?

OpenStudy (anonymous):

24 cm

OpenStudy (imstuck):

24 cm is the diameter. What is the radius?

OpenStudy (anonymous):

Oh, 12 cm

OpenStudy (imstuck):

Right! So can you finish the equation for the bull's eye of the target? x^2 + y^2 = what?

OpenStudy (anonymous):

it's r^2 right? Which is the same as the diameter, which is 24 cm

OpenStudy (imstuck):

No it's not the same. The radius is half the diameter. 12 * 2 = 24. 12^2 is a completely different thing. That's the same as 12 * 12. What's 12 * 12?

OpenStudy (anonymous):

12*12=144

OpenStudy (imstuck):

Right! So the equation for the smallest circle is x^2 + y^2 = 144

OpenStudy (imstuck):

Now onto the circle the next size up!

OpenStudy (imstuck):



OpenStudy (imstuck):

That's the bull's eye and the next size up circle. It has a width of 12, which I drew in, but how big is the radius now? Keep in mind that the smaller circle is now a part of the bigger circle. What is the radius of both circles together?

OpenStudy (anonymous):

24?

OpenStudy (imstuck):

That's right! Can you write the equation for that circle? Remember that it's still centered at the origin, and the bigger circle's radius is 24.

OpenStudy (anonymous):

Can we just use the old equation of x^2+y^2=144 and make it x^2+y^2=144^2?

OpenStudy (imstuck):

No because 144 squared is 20,736, and the radius of your circle is 24, and 24 squared is 576. See why you can't do that?

OpenStudy (anonymous):

Yeah, I see it now. So what do I do?

OpenStudy (imstuck):

The equation for the next size up circle is still the x^2 + y^2, but your radius squared is 576. So the equation is x^2 + y^2 = 576.

OpenStudy (imstuck):

Onto the next size up circle! There are 4 circles around the bull's eye and this one now is the second one.

OpenStudy (anonymous):

Is the equation for the next circle x^2+y^2=1296?

OpenStudy (imstuck):

Yes! What about the third one? The radius will be 12 * 4 which is 48. What is the equation?

OpenStudy (anonymous):

2304

OpenStudy (imstuck):

ok....one more to go. 12*5 = 60. Whats the equation?

OpenStudy (anonymous):

x^2+y^2=3600

OpenStudy (imstuck):

You got it! Good job! You got all of them now!

OpenStudy (anonymous):

Thanks! Can you help with another problem?

OpenStudy (imstuck):

sure!

OpenStudy (anonymous):

I'll make a new post

OpenStudy (anonymous):

Hey @IMStuck how do I draw this on a coordinate plane?